# arXiv:2603.02685v1[cond-mat.supr-con] 03 Mar 2026

### v1

@@ Abstract
As an unconventional superconducting system capable of reaching 40K under ambient pressure, the La3Ni2O7 film superconductor has become a recent focal point in the field of superconductivity, demanding further theoretical exploration of possible pairing mechanisms. In this work, we employ the fluctuation exchange (FLEX) approximation to systematically analyze the superconducting properties of the two-orbital, two-site model for the La3Ni2O7 film in the weakly correlated regime, focusing on its dependence on hole-doped concentration. Through a more detailed examination of the Fermi surface topology, we find that when a $\delta$ pocket composed of the $d_{z^{2}}$ orbital emerges near the $\Gamma$-point, its nesting with the $\gamma$-pocket, along with the nesting between the $\alpha$- and $\beta$-pockets, leads to a mutual enhancement of $s_{\pm}$-wave pairing at the corresponding wave vector. Furthermore, we propose that this nesting-driven enhancement of spin-fluctuation-induced pairing may be a viable mechanism to enhance superconductivity.
@@
The mechanism of high-temperature superconductivity has long been a central research topic in condensed matter physics. In recent years, following the discovery of cuprate and iron-based superconductors, the emergence of the nickel-based superconductor family LDF2019 ; Wang327_Nature2023 has attracted wide attention. Among them, Ruddlesden-Popper (RP) phase nickel-based superconductors, exemplified by the La3Ni2O7 compound, can achieve a superconducting transition temperature of up to 80K under high pressure conditions Wang327_Nature2023 . During the past two years, significant progress has been made in both experimental characterization hou2023emergence ; Zhang_2024_HighTc ; Xie_2024_MagneticExchange ; Li_2024_ElectronicCorrelation ; Wang_2024_La2PrNi2O7_Nature ; dan_2024 ; wang2024pressure ; cai_2024 ; liu_2024 and theoretical studies YaoPRL2023 ; Yang_2023_sPlusMinus ; Liu_2023_sPlusMinus ; Yang_2023_ValenceBonds ; Qin_2023_HighTc ; zhang_2023 ; pupha_l2023 ; Christiansson_2023 ; pupha_l2023 ; liao2023electron ; liao_2024 ; zhang2024electronic ; zhang2023trends ; Tian_2024_Correlation ; Oh_2024_StrongPairing ; Fan_2024_BilayerCoupling ; RyeePRL2024CMDFT ; wang2024electronic ; chen2024electronic ; lu2024interlayer ; sakakibara2024 ; jiang2024high ; zhang2024strong ; Zheng_2025_TwoOrbital ; jiang2025theory ; duan2025orbital ; xi2025transition ; shi2025theoretical ; ouyang2025two ; maier2025interlayer ; wang2025electronic ; puphal2025superconductivity of high-pressure nickel-based superconductors. One of the central questions in the theoretical study of nickel-based superconductors is how to further enhance their superconductivity, enabling them to operate at lower pressures and higher temperatures. The shape of the Fermi surface plays a critical role in determining superconducting behavior Wang327_Nature2023 ; Yang_2023_sPlusMinus ; Liu_2023_sPlusMinus ; Yang_2023_ValenceBonds . Importantly, the Fermi surface is sensitive to variations in lattice constants, electron filling, and interaction strength wu2024ac ; jiang2024high ; huo2025modulation ; wang2025electronic . Theoretically, many research groups have explored various strategies such as chemical doping, applying pressures to tune the Fermi surface geometry and the density of states at the Fermi level, trying to enhance superconductivity wang2023observation ; wu2024ac ; rhodes2024structural ; geisler2024structural ; chen2025charge ; huo2025modulation .
@@
Recently, the research team led by Z.Y.Chen achieved a breakthrough by successfully growing a La2.85Pr0.15Ni2O7 thin-film superconductor on an SrLaAlO4 substrate under ambient pressure, with a superconducting transition temperature reaching about 40K Ko_LNO327AmbientSC ; Zhou_2024_AmbientSC , stimulating extensive experimental and theoretical discussions regarding their superconducting mechanism liu2025superconductivity ; XJZhou_ARPES_327 ; li2025photoemission . The theoretical study by the authors yue2025correlated reveal this complex heterostructure is a moderately correlated system far from doped Mott insulator regime. Our results suggested that the superconductivity in ambient pressure La3Ni2O7 films may originate from spin-fluctuation-induced $s_{\pm}$-wave pairing.
@@ III Results
The $\delta$ band in realistic thin-film of La3Ni2O7 ever reported tends to be away from the Fermi level. However, by slightly modifying the lattice structure through in-plane strain LHY2507 , doping or substrate substitution, the $\delta$ pocket could be ”brought back” to the Fermi surface in new material realizations. If achieved in the future experiment, the presence of the $\delta$ pocket and its enhancement mechanism proposed in this study could lead to a higher superconducting transition temperature in nickel-based superconductors.
Acknowledgements. — Changming Yue acknowledges support from the National Natural Science Foundation of China (12474231), Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401004). Jian-Jian Miao is supported by NSFC (No. 12404171), the Guangdong Project (Grant No. 2024QN11X176).

### v2

# Possible Enhancement of Superconductivity in Ambient-Pressure La3Ni2O7 Thin Film

@@ Abstract
As an unconventional superconducting system capable of reaching 60 K under ambient pressure, the La3Ni2O7 thin film superconductor has recently become a focal point in the field of superconductivity, calling for further theoretical exploration of its possible pairing mechanisms. In this work, we employ the fluctuation exchange (FLEX) approximation to systematically analyze the superconducting properties of a previously proposed two-site, two-orbital model for the La3Ni22O77 thin film in the weakly correlated regime, with particular emphasis on its dependence on hole doping. Through a more detailed examination of the Fermi-surface topology, we find that when a $\delta$ pocket composed of the $d_{z^{2}}$ antibonding orbital emerges near the $\Gamma$ point, its nesting with the $\gamma$ pocket, together with the nesting between the $\alpha$ and $\beta$ pockets, leads to a mutual enhancement of $s{\pm}$-wave pairing at the corresponding wave vector. Furthermore, we propose that this nesting-driven enhancement of spin-fluctuation-induced pairing may provide a viable mechanism for enhancing superconductivity.
@@
The mechanism of high-temperature superconductivity has long been a central research topic in condensed matter physics. In recent years, following the discovery of cuprate and iron-based superconductors, the emergence of the nickel-based superconductor family LDF2019 ; Wang327_Nature2023 has attracted wide attention. Among them, Ruddlesden-Popper (RP) phase nickel-based superconductors, exemplified by the La3Ni2O7 compound, can achieve a superconducting transition temperature of up to 80K under high pressure conditions Wang327_Nature2023 . During the past two years, significant progress has been made in both experimental characterization hou2023emergence ; Zhang_2024_HighTc ; Xie_2024_MagneticExchange ; Li_2024_ElectronicCorrelation ; Wang_2024_La2PrNi2O7_Nature ; dan_2024 ; wang2024pressure ; cai_2024 ; liu_2024 and theoretical studies YaoPRL2023 ; Yang_2023_sPlusMinus ; Liu_2023_sPlusMinus ; Yang_2023_ValenceBonds ; Qin_2023_HighTc ; zhang_2023 ; pupha_l2023 ; Christiansson_2023 ; pupha_l2023 ; liao2023electron ; liao_2024 ; zhang2024electronic ; zhang2023trends ; Tian_2024_Correlation ; Oh_2024_StrongPairing ; Fan_2024_BilayerCoupling ; RyeePRL2024CMDFT ; wang2024electronic ; chen2024electronic ; lu2024interlayer ; sakakibara2024 ; jiang2024high ; zhang2024strong ; Zheng_2025_TwoOrbital ; jiang2025theory ; duan2025orbital ; xi2025transition ; shi2025theoretical ; ouyang2025two ; maier2025interlayer ; wang2025electronic ; puphal2025superconductivity ; xia2025sensitive ; jiang2025dual of high-pressure nickel-based superconductors. One of the central questions in the theoretical study of nickel-based superconductors is how to further enhance their superconductivity, enabling them to operate at lower pressures and higher temperatures. The shape of the Fermi surface plays a critical role in determining superconducting behavior Wang327_Nature2023 ; Yang_2023_sPlusMinus ; Liu_2023_sPlusMinus ; Yang_2023_ValenceBonds . Importantly, the Fermi surface is sensitive to variations in lattice constants, electron filling, and interaction strength wu2024ac ; jiang2024high ; huo2025modulation ; wang2025electronic . Theoretically, many research groups have explored various strategies such as chemical doping, applying pressures to tune the Fermi surface geometry and the density of states at the Fermi level, trying to enhance superconductivity wang2023observation ; wu2024ac ; rhodes2024structural ; geisler2024structural ; chen2025charge ; huo2025modulation .
@@
Recently, the research team led by Z.Y.Chen achieved a breakthrough by successfully growing a La2.85Pr0.15Ni2O7 thin-film superconductor on an SrLaAlO4 substrate under ambient pressure, with a superconducting transition temperature reaching about 40K Ko_LNO327AmbientSC ; Zhou_2024_AmbientSC . Recently, by refining the growth techniques, they further increased the superconducting transition temperature to 60 K zhou2025superconductivity , stimulating extensive experimental and theoretical discussions regarding their superconducting mechanism liu2025superconductivity ; XJZhou_ARPES_327 ; li2025photoemission ; ryee2025superconductivity ; ushio2025theoretical ; shi2025effect ; yue2025correlated ; gao2025theoretical ; wang2026superconducting . In particular, the Fermi-surface topology has become a central issue in recent discussions of the superconducting mechanism ryee2025superconductivity ; ushio2025theoretical ; shi2025effect ; gao2025theoretical . The theoretical study by the authors yue2025correlated reveal this complex heterostructure is a moderately correlated system far from doped Mott insulator regime. Our results suggested that the superconductivity in ambient pressure La3Ni2O7 films may originate from spin-fluctuation-induced $s_{\pm}$-wave pairing.
@@ III Results
The $\delta$ band in realistic thin-film of La3Ni2O7 ever reported tends to be away from the Fermi level. However, by slightly modifying the lattice structure through in-plane strain LHY2507 , doping or substrate substitution, the $\delta$ pocket could be restored to the Fermi surface in new material realizations. If achieved in the future experiment, the presence of the $\delta$ pocket and its enhancement mechanism proposed in this study could lead to a higher superconducting transition temperature in nickel-based superconductors.
Acknowledgements. — Changming Yue acknowledges support from the Guangdong Major Project of Basic Research (2025B0303000004), Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401004, GDZX2501001), the National Natural Science Foundation of China (12474231). Jian-Jian Miao is supported by NSFC (No. 12404171), the Guangdong Project (Grant No. 2024QN11X176). Wei-Qiang Chen was supported by the National Key Research and Development Program of China (No. 2024YFA1408101), NSFC (Grants No. 12334002), the SUSTech-NUS Joint Research Program, the Science, Technology and Innovation Commission of Shenzhen Municipality (No. ZDSYS20190902092905285), and Center for Computational Science and Engineering at Southern University of Science and Technology.